# mystyle backgroundcolor=white!10

### v1

@@ Abstract
This paper studies finite-time safety and reach-avoid verification for stochastic discrete-time dynamical systems. The aim is to ascertain lower and upper bounds of the probability that, within a predefined finite-time horizon, a system starting from an initial state in a safe set will either exit the safe set (safety verification) or reach a target set while remaining within the safe set until the first encounter with the target (reach-avoid verification). We introduce novel barrier-like sufficient conditions for characterizing these bounds, which either complement existing ones or fill gaps. Finally, we demonstrate the efficacy of these conditions on two examples.
Index Terms:
Finite-time Safety/Reach-avoid Verification; Probability Bounds; Stochastic Discrete-time Systems
I Introduction
@@
Barrier certificates were initially proposed for deterministic systems as a popular formal approach to temporal verification in [8]. They offer Lyapunov-like assurances regarding system behavior, with the mere existence of a barrier function being sufficient to establish the satisfiability of safety and/or reachability specifications in [10]. Subsequent efforts have focused on adapting and enhancing deterministic barrier functions, as well as broadening their applications [2, 14]. However, many real-world applications are susceptible to stochastic disturbances and are thus modeled as stochastic systems. In the stochastic setting, safety verification over the infinite time horizon via barrier certificates was introduced alongside its deterministic counterpart in [9]. Utilizing Ville’s Inequality [12], [9] constructed a non-negative barrier function and provided a sufficient condition for upper bounding the probability of eventually entering an unsafe region from a given initial state while remaining within a state constraint set. More recently, a new barrier function, constructed by relaxing a set of equations, was proposed for lower bounding the probability of eventually entering an unsafe or desired region from an initial state while adhering to state constraints in [13, 15]. These barrier functions were further extended to determine the lower and upper bounds of the liveness probability over the infinite time horizon for a specified safe set and set of initial states in [16]. Notably, these proposed barrier functions require the expected value evolution to be monotonically non-increasing or non-decreasing, which can be demanding and thus limit their practical applications. On the other hand, finite-time temporal verification holds greater practical significance as most real-world systems operate within well-defined time constraints. Drawing inspiration from [6], the concept of a $c$-martingale was introduced in [11] for stochastic continuous-time systems modelled by stochastic differential equations, enabling a controlled increase in the expected certificate value at each time step and offering an upper bound on the probability of reaching unsafe sets within bounded time horizons. Building upon this, [5] extended the $c$-martingale framework to address temporal logic verification for discrete-time systems, further extending the applicability of this approach.
@@
This paper investigates the finite-time safety and reach-avoid verification of stochastic discrete-time systems. The finite-time safety verification problem aims to compute both lower and upper bounds of the probability that a system, starting from an initial state in a safe set, will exit the safe set throughout a given bounded time interval. From a reachability perspective, it involves computing lower and upper bounds on the probability of reaching the complement of the safe set within the specified bounded time interval. Thus, it exclusively addresses safety or reachability concerns. In contrast, finite-time reach-avoid verification integrates guarantees of safety and reachability. It seeks to establish lower and upper bounds on the probability that a system, starting from an initial state in a safe set, will reach a target set within a designated bounded time interval while ensuring it remains within the safe set before reaching the target set. Although these two problems are interconnected, they differ fundamentally in essence, as we will elaborate on in the preliminaries section. We propose novel barrier-like conditions to address these two problems. These conditions either complement existing ones or fill gaps, facilitating the attainment of tight probability bounds. Finally, we demonstrate the effectiveness of the proposed conditions on two numerical examples, utilizing semi-definite programming tools.
@@
    This work studies the finite-time safety and reach-avoid verification of stochastic discrete-time systems. Compared with existing works [5, 6], which merely offered upper bounds on the relevant probabilities, our contribution goes beyond by providing both lower and upper bounds. These bounds deepen our understanding of the system’s characteristics and yield a more accurate estimation of the true probability, thus enhancing the overall rigour and precision of the finite-time safety and reach-avoid analysis.
@@
    The proposed barrier-like conditions for upper bounding the probabilities complement existing ones in [5, 6], facilitating the gain of sharper upper bounds.
This paper is structured as follows: in Section II, we formulate the finite-time safety and reach-avoid verification problems of interest. In Section III and IV, we introduce our barrier-like conditions for addressing these two problems, respectively. After demonstrating the performance of the proposed conditions on two examples in Section V, we conclude this paper in Section VI.
II Preliminaries
We start the exposition by a formal introduction of discrete-time systems subject to stochastic disturbances and finite-time safety/reach-avoid verification problems of interest. Before posing the problem studied, let me introduce some basic notions used throughout this paper: $\mathbb{R}$ denotes the set of real values; $\mathbb{N}$ denotes the set of nonnegative integers; $\mathbb{N}_{\leq k}$ is the set of non-negative integers being less than or equal to $k$; for sets $\Delta_{1}$ and $\Delta_{2}$, $\Delta_{1}\setminus\Delta_{2}$ denotes the difference of sets $\Delta_{1}$ and $\Delta_{2}$, which is the set of all elements in $\Delta_{1}$ that are not in $\Delta_{2}$; $1_{A}(\bm{x})$ denotes the indicator function in the set $A$, where, if $\bm{x}\in A$, then $11_{A}(\bm{x})=$ 1 and if $\bm{x}\notin A$, $01_{A}(\bm{x})=$ 0.
II-A Problem Statement
@@
where $\bm{x}(\cdot)\colon\mathbb{N}\rightarrow\mathbb{R}^{n}$ are states and $\bm{\theta}(\cdot)\colon\mathbb{N}\rightarrow\Theta$ with $\Theta\subseteq\mathbb{R}^{m}$ are stochastic disturbances. In addition, let $\bm{\theta}(0),\bm{\theta}(1),\ldots$ be i.i.d. random vectors distributed according to $\mathbb{P}$, with support $\Theta$. The expectation $\mathbb{E}[\cdot]$ is defined with respect to this distribution.
Before defining the trajectory of system (1), we define a disturbance signal.
Definition 1
A disturbance signal $\pi$ is an ordered sequence $\{\bm{\theta}(i),i\in\mathbb{N}\}$, where $\bm{\theta}(\cdot)\colon\mathbb{N}\rightarrow\Theta$.
Given system (1), a signal $\pi=\{\bm{\theta}(i),i\in\mathbb{N}\}$ is a stochastic process defined on the canonical sample space $\Omega=\Theta^{\infty}$, endowed with its product topology $\mathcal{B}(\Omega)$, with the probability measure $\mathbb{P}^{\infty}$. The expectation associated with the probability measure $\mathbb{P}^{\infty}$ is denoted by $\mathbb{E}^{\infty}[\cdot]$.
@@
Definition 2
Given a disturbance signal $\pi\in\Omega$ and an initial state $\bm{x}_{0}\in\mathbb{R}^{n}$, a trajectory of system (1) is denoted as $\bm{\phi}_{\pi}^{\bm{x}_{0}}(\cdot)\colon\mathbb{N}\rightarrow\mathbb{R}^{n}$ with $\bm{\phi}_{\pi}^{\bm{x}_{0}}(0)=\bm{x}_{0}$, i.e., $\bm{\phi}_{\pi}^{\bm{x}_{0}}(l+1)=\bm{f}(\bm{\phi}_{\pi}^{\bm{x}_{0}}(l),\bm{\theta}(l)),\forall l\in\mathbb{N}.$
In this study, we address two verification problems for the system governed by (1) with a finite-time horizon $[0,N]$, where $N\in\mathbb{N}$. The first problem pertains to a finite-time safety verification problem, examining the likelihood of the system exiting the safe set $\mathcal{X}$ throughout its evolution over $[0,N]$, starting from $\bm{x}_{0}\subseteq\mathcal{X}$. The second one involves a finite-time reach-avoid task, focusing on the probability that the system enters the target set $\mathcal{X}_{r}\subset\mathcal{X}$ safely within the time horizon $[0,N]$, given an initial state $\bm{x}_{0}\in\mathcal{X}\setminus\mathcal{X}_{r}$.
Problem 1 (Finite-time Safety Verification)
@@
Problem 2 (Finite-time Reach-avoid Verification)
Given a finite time interval $[0,N]$ with $N\in\mathbb{N}$, a safe set $\mathcal{X}$, a target set $\mathcal{X}_{r}\subseteq\mathcal{X}$, and an initial state $\bm{x}_{0}\in\mathcal{X}\setminus\mathcal{X}_{r}$, the finite-time reach-avoid verification problem is to compute $\epsilon_{1}\in[0,1]$ and $\epsilon_{2}\in[0,1]$ which are respectively the lower and upper bounds of the probability, with which system (1) starting from $\bm{x}_{0}$ will enter the target set $\mathcal{X}_{r}$ within $[0,N]$ while staying inside the set $\mathcal{X}$ before the first target hitting time, i.e.,
$\epsilon_{1}\leq\mathbb{P}^{\infty}\left(\begin{aligned} &\exists k\in\mathbb{N}_{\leq N}.\bm{\phi}_{\pi}^{\bm{x}_{0}}(k)\in\mathcal{X}_{r}\wedge\forall l\in\mathbb{N}_{\leq k}.\bm{\phi}_{\pi}^{\bm{x}_{0}}(l)\in\mathcal{X}\end{aligned}\middle|\bm{x}_{0}\in\mathcal{X}\setminus\mathcal{X}_{r}\right)\leq\epsilon_{2}.$
@@
III Finite-time Safety Verification
In this section, we present our sufficient conditions for characterizing upper and lower bounds of the probability in Problem 1. The sufficient condition for lower bounds is formulated in Subsection III-A and the one for upper bounds is introduced in Subsection III-B.
@@ III Finite-time Safety Verification
Definition 3
The switched stochastic discrete-time system, which is built upon system (1), is a quadruple (~$(\widetilde{\mathcal{L}},\widetilde{\mathcal{X}},\bm{x}_{0},\widetilde{\bm{f}})$ with the following components:
@@
    ~=$\widetilde{\mathcal{L}}=\{1,2\}$ is a set of three locations;
@@
    with $\widetilde{\bm{f}}_{1}(\bm{x},\bm{\theta})=\bm{f}(\bm{x},\bm{\theta})$ and $\widetilde{\bm{f}}_{2}(\bm{x},\bm{\theta})=\bm{x}$, and $1_{\widetilde{\mathcal{X}}_{i}}(\bm{x})$ is the indicator function of the set $\widetilde{\mathcal{X}}_{i}$, i.e., $11_{\widetilde{\mathcal{X}}_{i}}(\bm{x})=$ 1 if $\bm{x}\in\widetilde{\mathcal{X}}_{i}$; otherwise, $01_{\widetilde{\mathcal{X}}_{i}}(\bm{x})=$ 0,
@@
    $\widetilde{\Omega}=\{\bm{x}\in\mathbb{R}^{n}\mid\bm{x}=\bm{f}(\bm{x}_{0},\bm{\theta}),\bm{x}_{0}\in\mathcal{X},\bm{\theta}\in\Theta\}\cup\mathcal{X};$
@@
$\begin{split}&\bm{x}(l+1)=\widetilde{\bm{f}}(\bm{x}(l),\bm{\theta}(l)),\forall l\in\mathbb{N},\\ &\bm{x}(0)=\bm{x}_{0}\in\widetilde{\mathcal{X}}.\end{split}$ (2)
The trajectory of system (2), induced by initial state $\bm{x}_{0}\in\widetilde{\mathcal{X}}$ and disturbance policy $\pi\in\Omega$, is denoted by $\widetilde{\bm{\phi}}_{\pi}^{\bm{x}_{0}}(\cdot)\colon\mathbb{N}\rightarrow\mathbb{R}^{n}$.
@@
III-A Sufficient Conditions for Upper Bounds
In this subsection, we present our sufficient condition for upper bounding the probability of exiting the safe set over the time horizon $[0,N]$ in Problem 1.
Theorem 1

### v2

# Finite-time Safety and Reach-avoid Verification of Stochastic Discrete-time Systems

@@ Abstract
This paper studies finite-time safety and reach-avoid verification for stochastic discrete-time dynamical systems. The aim is to ascertain lower and upper bounds of the probability that, within a predefined finite-time horizon, a system starting from an initial state in a safe set will either exit the safe set (safety verification) or reach a target set while remaining within the safe set until the first encounter with the target (reach-avoid verification). We introduce novel barrier-like sufficient conditions for characterizing these bounds, which either complement existing ones or fill gaps. Finally, we demonstrate the efficacy of these conditions on two simple examples.
keywords:
Stochastic Discrete-time Systems , Finite-time Safety Verification , Finite-time Reach-avoid Verification , Barrier-like Conditions , Lower and Upper Probability Bounds
[…]
organization=Key Laboratory of System Software (Chinese Academy of Sciences) and State Key Laboratory of Computer Science, Institute of Software, Chinese Academy of Sciences, city=Beijing, China, postcode=100190, country=China
Email: xuebai@ios.ac.cn
1 Introduction
@@
Barrier certificates were initially proposed for deterministic systems as a popular formal approach to temporal verification in [23]. They offer Lyapunov-like assurances regarding system behavior, with the mere existence of a barrier function being sufficient to establish the satisfiability of safety and/or reachability specifications in [25]. Subsequent efforts have focused on adapting and enhancing deterministic barrier functions, as well as broadening their applications [4, 29, 35]. However, many real-world applications are susceptible to stochastic disturbances and are thus modeled as stochastic systems. In the stochastic setting, safety verification over the infinite time horizon via barrier certificates was introduced alongside its deterministic counterpart in [24]. Utilizing Ville’s Inequality [30], [24] constructed a non-negative barrier function and provided a sufficient condition for upper bounding the probability of eventually entering an unsafe region from a given initial state while remaining within a state constraint set. More recently, a new barrier function, constructed by relaxing a set of equations, was proposed for lower bounding the probability of eventually entering an unsafe or desired region from an initial state while adhering to state constraints in [34, 36]. These barrier functions were further extended to determine the lower and upper bounds of the safety probability over the infinite time horizon for a specified safe set and set of initial states in [37]. Furthermore, by formulating barrier certificates via a relaxation of Bellman equations, [33] established necessary and sufficient conditions for lower and upper bounds the safety and reach-avoid probabilities of stochastic discrete-time systems over the infinite time horizon. In addition, under the assumption that a robust invariant set exists and the system evolves within this robust invariant set, [39] proposed a new barrier certificate, termed reach-avoid supermartingale, to guarantee satisfaction of reach-avoid specifications as well as facilitate reach-avoid controllers. Barrier certificates have also been extended to infinite-time horizon probabilistic program analysis, where they verify properties such as (positive) almost-sure termination, probabilistic termination, assertion violations, and reachability (e.g., [6, 21, 22, 15, 9, 7, 18, 32, 28]). Specially, the probabilistic termination analysis discussed in [9, 7, 18] exhibits interesting connections with classical analysis for stochastic discrete-time systems. When seeking a lower bound for termination probability, this analysis shares conceptual similarities with classical reach-avoid analysis. This connection arises because the termination probability essentially measures the likelihood that a program, beginning from a specific initial state, will eventually reach terminal states while preserving a stochastic invariant set throughout its execution prior to termination. Conversely, the probabilistic termination analysis presented in [18], which focuses on establishing upper bounds for termination probability, parallels classical safety analysis in stochastic discrete-time systems. In this case, the termination probability represents the chance that a program will eventually violate (escape from) a stochastic invariant set (which excludes the terminal state) when starting from a given initial state. Moreover, the application of barrier certificates has been expanded to encompass both qualitative and quantitative analysis of $\omega$-regular properties [11, 1, 2, 12].
On the other hand, finite-time temporal verification holds greater practical significance as most real-world systems operate within well-defined time constraints. Drawing inspiration from [17], the concept of a $c$-martingale was introduced in [27] for stochastic continuous-time systems modelled by stochastic differential equations, enabling a controlled increase in the expected certificate value at each time step and offering an upper bound on the exit probability of leaving safe sets within bounded time horizons. Afterwards, [20] proposed a computational method to find a $c$-martingale expressed by neural networks for finite-time safety verification of stochastic discrete-time systems. The $c$-martingales in [20] are a typical case of the proposed barrier function in the present work. Under the assumption that the system evolves within a robust invariant set, [13] extended the $c$-martingale framework to address temporal logic verification for discrete-time systems, and later, [14] synthesized control policies for discrete-time stochastic control systems together with a lower bound on the probability that the systems satisfy complex temporal properties. [26] utilized barrier-like results introduced in [17] to address the challenges of finite-time safety verification and the synthesis of safe controllers for stochastic discrete-time systems, employing semi-definite programming techniques. The detailed description of the verification problem in [26] and its relationship with the verification problem are illustrated in Remark 3. On the other hand, the aforementioned works on finite-time temporal verification offer only upper bounds of the probability of reaching undesirable sets (equivalently, lower bounds of the probability of staying within desirable sets). Such works do not provide the lower bounds. This work will offer both types of bounds. Recent work by [38] introduced barrier functions for bounding probabilistic safety (lower and upper bounds) across infinite and finite time horizons. Its analysis, similar to [13, 14], relies on the assumption that system evolution remains within a robust invariant set. However, as critically examined in [37], this strong invariance requirement constitutes a significant limitation. In contrast, the current work develops new barrier functions that eliminate the need for such a restrictive assumption. Similar to the infinite-time case, barrier certificates have also been extended to the analysis of probabilistic programs within bounded time horizons (e.g., [9, 16, 8]). [9] (e.g., Lemma 3) presented a sufficient condition based on an $\epsilon$-repulsing supermartingale supported by a pure invariant, utilizing Azuma’s and Hoeffding’s inequalities to derive an upper bound for programs that reach a specified set exactly at a given step. Afterward, in the context where programs terminate almost surely, [16] proposed a sufficient condition for establishing lower bounds on program termination within bounded time horizons; [31] developed lower and upper bounds for the tail bound problem which can also be employed to bound the probability that programs terminate within bounded time horizons. Recently, in conjunction with stochastic invariants [9], [8] investigated the tail bound problem for programs that do not necessarily terminate almost surely. This approach can also be employed to establish lower bounds on the probability that programs terminate within bounded time horizons.
@@
This paper investigates the finite-time safety and reach-avoid verification of stochastic discrete-time systems. The finite-time safety verification problem aims to compute both lower and upper bounds of the probability that a system, starting from an initial state in a safe set, will exit the safe set throughout a given bounded time interval. From a reachability perspective, it involves computing lower and upper bounds on the probability of reaching the complement of the safe set within the specified bounded time interval. Thus, it exclusively addresses safety or reachability concerns. In contrast, finite-time reach-avoid verification integrates guarantees of safety and reachability. It seeks to establish lower and upper bounds on the probability that a system, starting from an initial state in a safe set, will reach a target set within a designated bounded time interval while ensuring it remains within the safe set before reaching the target set. Although these two problems are interconnected, they differ fundamentally in essence, as we will elaborate on in the preliminaries section. We propose novel barrier-like conditions to address these two problems. These conditions either complement existing ones or fill gaps, facilitating the attainment of tight probability bounds for some systems. Finally, we demonstrate the effectiveness of the proposed conditions on two numerical examples, utilizing semi-definite programming tools.
@@
    This work studies the finite-time safety and reach-avoid verification of stochastic discrete-time systems. Compared with existing works [26, 20, 17], which merely offered upper bounds on the relevant probabilities, our contribution goes beyond by providing both lower and upper bounds. These bounds deepen our understanding of the system’s characteristics and yield a more accurate estimation of the true probability, thus enhancing the overall rigour and precision of the finite-time safety and reach-avoid analysis. In addition, obtaining both lower and upper bounds also facilitates evaluating their mutual tightness in practice.
@@
    The proposed barrier-like conditions for upper bounding the probabilities complement existing ones in [20, 17], facilitating the gain of sharper upper bounds for some systems.
  3. 3.
    Unlike the approach in [38], our proposed barrier-like conditions for bounding probabilities in finite-time safety verification eliminate the need for the strong invariance assumption. Additionally, this work extends these barrier-like conditions to finite-time reach-avoid verification, providing both lower and upper probability bounds.
This paper is structured as follows: in Section 2, we formulate the finite-time safety and reach-avoid verification problems. In Sections 3 and 4, we introduce our barrier-like conditions for addressing these two problems, respectively. After demonstrating the performance of proposed conditions on two examples in Section 5, we conclude this paper in Section 6.
2 Preliminaries
We start the exposition by a formal introduction of discrete-time systems subject to stochastic disturbances and finite-time safety/reach-avoid verification problems of interest. Before posing the problem studied, let me introduce some basic notions used throughout this paper: $\mathbb{R}$ denotes the set of real values; $\mathbb{N}$ denotes the set of nonnegative integers; $\mathbb{N}_{\leq k}$ is the set of non-negative integers being less than or equal to $k$; for sets $\Delta_{1}$ and $\Delta_{2}$, $\Delta_{1}\setminus\Delta_{2}$ denotes the difference of sets $\Delta_{1}$ and $\Delta_{2}$, which is the set of all elements in $\Delta_{1}$ that are not in $\Delta_{2}$; $1_{A}(\bm{x})$ denotes the indicator function in the set $A$, where, if $\bm{x}\in A$, then $1_{A}(\bm{x})=1$ and if $\bm{x}\notin A$, $1_{A}(\bm{x})=0$.
@@ 2 Preliminaries
where $\bm{x}(l)\in\mathbb{R}^{n}$ is the state at time $l$ and $\bm{\theta}(l)\in\Theta$ with $\Theta\subseteq\mathbb{R}^{m}$ is the stochastic disturbance at time $l$. In addition, let $\bm{\theta}(0),\bm{\theta}(1),\ldots$ be i.i.d. (independent and identically distributed) random vectors on a probability space $(\Theta,\mathcal{F},\mathbb{P})$, and take values in $\Theta$ with the following probability distribution: for any measurable set $B\subseteq\Theta$,
${\rm Prob}(\bm{\theta}(l)\in B)=\mathbb{P}(B),\quad\forall l\in\mathbb{N}.$
The expectation $\mathbb{E}[\cdot]$ is defined with respect to the probability measure $\mathbb{P}$.
Before defining trajectories of system (1), we define a disturbance signal.
Definition 1.
A disturbance signal $\pi$ is a sample path of the stochastic process $\{\bm{\theta}(i)\colon\Theta\rightarrow\Theta,i\in\mathbb{N}\}$, which is defined on the canonical sample space $\Theta^{\infty}$, endowed with its product topology $\mathcal{B}(\Theta^{\infty})$, with the probability measure $\mathbb{P}^{\infty}$. The expectation associated with the probability measure $\mathbb{P}^{\infty}$ is denoted by $\mathbb{E}^{\infty}[\cdot]$.
@@
Definition 2.
Given a disturbance signal $\pi$ and an initial state $\bm{x}_{0}\in\mathbb{R}^{n}$, a trajectory of system (1) is denoted as $\bm{\phi}_{\pi}^{\bm{x}_{0}}(\cdot)\colon\mathbb{N}\rightarrow\mathbb{R}^{n}$ with $\bm{\phi}_{\pi}^{\bm{x}_{0}}(0)=\bm{x}_{0}$, i.e., $\bm{\phi}_{\pi}^{\bm{x}_{0}}(l+1)=\bm{f}(\bm{\phi}_{\pi}^{\bm{x}_{0}}(l),\bm{\theta}(l)),\forall l\in\mathbb{N}.$
In this study, we address two verification problems for the system governed by (1) with a finite time horizon $[0,N]$, where $N\in\mathbb{N}$. The first problem pertains to a finite-time safety verification problem, examining the likelihood of the system exiting the safe set $\mathcal{X}$ throughout its evolution over $[0,N]$, starting from $\bm{x}_{0}\subseteq\mathcal{X}$. The second one involves a finite-time reach-avoid task, focusing on the probability that the system enters the target set $\mathcal{X}_{r}\subset\mathcal{X}$ safely within the time horizon $[0,N]$, given an initial state $\bm{x}_{0}\in\mathcal{X}\setminus\mathcal{X}_{r}$.
Problem 1 (Finite-time Safety Verification).
@@
Problem 2 (Finite-time Reach-avoid Verification).
Given a finite time interval $[0,N]$ with $N\in\mathbb{N}$, a safe set $\mathcal{X}$, a target set $\mathcal{X}_{r}\subseteq\mathcal{X}$, and an initial state $\bm{x}_{0}\in\mathcal{X}\setminus\mathcal{X}_{r}$, the finite-time reach-avoid verification problem is to compute $\epsilon_{1}\in[0,1]$ and $\epsilon_{2}\in[0,1]$ which are respectively the lower and upper bounds of the reach-avoid probability, with which system (1) starting from $\bm{x}_{0}$ will enter the target set $\mathcal{X}_{r}$ within $[0,N]$ while staying inside the set $\mathcal{X}$ before the first target hitting time, i.e.,
$\epsilon_{1}\leq\mathbb{P}^{\infty}\left(\begin{aligned} &\exists k\in\mathbb{N}_{\leq N}.\bm{\phi}_{\pi}^{\bm{x}_{0}}(k)\in\mathcal{X}_{r}\\ &\wedge\forall l\in\mathbb{N}_{\leq k}.\bm{\phi}_{\pi}^{\bm{x}_{0}}(l)\in\mathcal{X}\end{aligned}\middle|\bm{x}_{0}\in\mathcal{X}\setminus\mathcal{X}_{r}\right)\leq\epsilon_{2}.$
@@
3 Finite-time Safety Verification
In this section, we present our sufficient conditions for characterizing upper and lower bounds on the probability in Problem 1. The sufficient condition for lower bounds is formulated in Subsection 3.1 and the one for upper bounds is introduced in Subsection 3.2.
@@ 3 Finite-time Safety Verification
Definition 3.
[…]
@@
    $\widetilde{\mathcal{L}}=\{1,2\}$ is a set of two locations;
@@
    with $\widetilde{\bm{f}}_{1}(\bm{x},\bm{\theta})=\bm{f}(\bm{x},\bm{\theta})$ and $\widetilde{\bm{f}}_{2}(\bm{x},\bm{\theta})=\bm{x}$, and $1_{\widetilde{\mathcal{X}}_{i}}(\bm{x})$ is the indicator function of the set $\widetilde{\mathcal{X}}_{i}$, i.e., $1_{\widetilde{\mathcal{X}}_{i}}(\bm{x})=1$ if $\bm{x}\in\widetilde{\mathcal{X}}_{i}$; otherwise, $1_{\widetilde{\mathcal{X}}_{i}}(\bm{x})=0$,
@@
    $\widetilde{\Omega}=\{\bm{x}^{\prime}\in\mathbb{R}^{n}\mid\bm{x}^{\prime}=\bm{f}(\bm{x},\bm{\theta}),\bm{x}\in\mathcal{X},\bm{\theta}\in\Theta\}\cup\mathcal{X};$
@@
[…]
The trajectory of system (2), induced by initial state $\bm{x}_{0}\in\widetilde{\mathcal{X}}$ and disturbance policy $\pi$, is denoted by $\widetilde{\bm{\phi}}_{\pi}^{\bm{x}_{0}}(\cdot)\colon\mathbb{N}\rightarrow\mathbb{R}^{n}$.
@@
3.1 Sufficient Conditions for Upper Bounds
In this subsection, we present our sufficient condition for upper bounding the probability of exiting the safe set $\mathcal{X}$ over the time horizon $[0,N]$ in Problem 1.
The sufficient condition requires the existence of a non-negative function $v(\bm{x}):\widetilde{\mathcal{X}}\rightarrow\mathbb{R}$ defined on the set $\widetilde{\mathcal{X}}$ that satisfies two key properties: (1) it must be greater than or equal to one on the subset $\widetilde{\mathcal{X}}\setminus\mathcal{X}$, and (2) its $\alpha$-scaled expected value at the next step along the system dynamics described in (1) should not increase by more than $\alpha\beta$ relative to its current value for $\bm{x}\in\mathcal{X}$, where $\alpha\in(0,1]$ and $\beta\in(-\infty,1]$. The upper bounds derivation proceeds as follows: Lemma 1 establishes that the probability of system (2) entering $\widetilde{\mathcal{X}}\setminus\mathcal{X}$ at time $t=N$ is equal to the probability of system (1) entering this set up to time $t=N$. Then, we derive upper bounds using system (2) by analyzing $\alpha$ and $\beta$. Relying on the fact that the system in (2) remains stationary when starting from $\widetilde{\mathcal{X}}\setminus\mathcal{X}$, which implies $\mathbb{E}^{\infty}[v(\widetilde{\bm{\phi}}_{\pi}^{\bm{x}}(1))]=v(\bm{x})$ for $\bm{x}\in\widetilde{\mathcal{X}}\setminus\mathcal{X}$, we reformulate the constraint over the function $v(\bm{x}):\widetilde{\mathcal{X}}\rightarrow\mathbb{R}$ using system (2), requiring that the $\alpha$-scaled expected value of $v(\bm{x})$ at the next step along the system (2) grows by at most $\alpha\beta$ relative to its current value $v(\bm{x})$ over the whole set $\widetilde{\mathcal{X}}$ when $\alpha\in(0,1]$ and $\beta\alpha-(\alpha-1)\in[0,1]$. Upper bounds follow from recursive application of this reformulated constraint; when $\alpha\in(0,1]$ and $\beta\alpha-(\alpha-1)\in(-\infty,0)$, the constraint is modified to require that the expected value of $v(\bm{x})$ at the next step along the system (2) grows by at most $\beta$, relative to the $(1-\beta)$-weighted current value $v(\bm{x})$. The upper bound again emerges through recursive application of this reformulated constraint.
Theorem 1.